OpenStudy (bebe):

find the difference enter you answer in simplest terms using the slash bar (/) as the fraction bar .5/6 - 1/8

OpenStudy (anonymous):

.5 = 1/2 (1/2)/(6/1) = (1/2)(1/6) = (1/12) (1/12) - (1/8) = -1/24 -1/24